OpenStudy (anonymous):

what is the slope of the line through (1,9) and (5,3)

OpenStudy (lgbasallote):

\[\large \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (lgbasallote):

\[\large m = \frac{3 - 9}{5-1}\]

OpenStudy (lgbasallote):

\[\large m = \frac{-6}{4}\]

OpenStudy (lgbasallote):

\[\large m = -\frac{3}{2}\]

OpenStudy (apoorvk):

Slope 'm' of a line passing through to points (x',y') and (x",y") is: \[m=\frac{y'-y"}{x'-x"}\] You have your two points inhere - just substitute and calculate.